OpenStudy (thomasfire55):

Solve the following quadratic equation by the square root property of equality. x^2 - 81 = 0

OpenStudy (prathamesh_m):

Try using: \[a ^{2}-b ^{2}=(a+b)(a-b)\]

OpenStudy (prathamesh_m):

The given equation is in this form.

OpenStudy (thomasfire55):

x=9?

jhonyy9 (jhonyy9):

x^2 -81 = 0 (x-9)(x+9) = 0 x_1 = 9 x_2 = - 9

jhonyy9 (jhonyy9):

but you can writing x^2 -81 = 0 x^2 = 81 x^2 = 9^2 x_1,2 = +/- 9